MEASUREMENT OF NITROGEN HYPERFINE STRUCTURE ON THE 53 CM (562 MHz) BUTYRONI-TRILE LINE

\underline{CHRISTOPHER T. DEWBERRY}, GARRY S. GRUBBS II, ANDREW RAPHELT^{a} and STEPHEN A. COOKE, The Department of Chemistry, University of North Texas, P.O. Box 305070, Denton, TX, 76203-5070.

Recent improvements to our cavity-based Fourier transform radiofrequency spectrometer will be presented. Amongst other improvements use of Miteq amp, model AMF-6F-00100400-10-10P (0.1 GHz to 4 GHz, 65 dB gain minimum, 1 dB noise figure maximum) together with shielding from an improved Faraday cage have significantly helped us in this regard. Electromagnetic fields within our near-spherical cavity have been modeled and results will be presented. We have been able to easily resolve the nitrogen hyperfine structure on the ^{a}Q_{0,−1} transition 1_{1,0} ← 1_{1,1} located at 562 MHz. This result will be discussed.

aNSF-REU student from Concordia University, Nebraska